Update on work in HEAD
dillon at apollo.backplane.com
Fri Jun 10 15:08:35 PDT 2005
HEAD should be reasonably stable again. There are still a few possible
issues on SMP boxes so I am particularly interested in testing on those.
There are still a number of open bug reports but except for an issue
that David Rhodus and I have been trying to track down I've lost track
of what might still be open. If you are still having problems with
recent work please post an update describing the issue to the bugs@ list.
The SPL->CRIT work is in the final stretch! There are only 117 files
left to go, out of the original 500 or so. Committers who have time
should continue to request chunks until I've announced that all files
have been assigned!
<dillon at xxxxxxxxxxxxx>
More information about the Kernel